Pure copper was mixed with a 12% copper alloy to produce an alloy that was 32% copper. How much of the pure copper and how much 12% alloy were used to produce 44 Kg of 32% alloy?
Answer by mananth(16946)

percent ---------------- quantity
Pure Copper 100.00% ---------------- x kg
Copper Alloy 12.00% ------ 44 - x kg
Mixture 32.00% ---------------- 44
12
100.00% x+12.00% (44-x)=32.00%*44
100x+12(44-x)=1408
100x+528 -12x=1408
100x-12 x =1408-528
88x= 880
/ 88
x=10kg 100.00% Pure Copper
34 kg 12.00% Copper Alloy

m.ananth@hotmail.ca